OpenStudy (anonymous):

Solve the triangle. A = 33°, a = 20, b = 13

OpenStudy (anonymous):

Wat u Have to Find ?

OpenStudy (anonymous):

The options are : Cannot be solved B = 20.7°, C = 146.3°, c ≈ 23.7 B = 20.7°, C = 126.3°, c ≈ 17.8 B = 20.7°, C = 126.3°, c ≈ 29.6 @yahoo!

OpenStudy (anonymous):

Use Sine Law

OpenStudy (anonymous):

a/SinA = b/SinB

OpenStudy (anonymous):

From here Find Angle B then To Find Angle C Use This Relation A + B + C = 180

OpenStudy (anonymous):

wouldnt that make it option 2 ? @yahoo!

OpenStudy (anonymous):

Is side a next to angle A or opposite?

OpenStudy (anonymous):

 This is how you find B. Then to find C you plug A and B into this equation: A+B+C=180. The from there you use SinA/a = SinC/c to get c. If you need any more help tag me